OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 7, an = an-1 + 6

OpenStudy (anonymous):

13, 19, 25, 31, 37, 43 0, 6, 12, 18, 24, 30 7, 6, 12, 18, 24, 30 7, 13, 19, 25, 31, 37

OpenStudy (anonymous):

@nincompoop @Luigi0210

OpenStudy (anonymous):

just plug-in the values 1...6 for n...

OpenStudy (anonymous):

so its the third choice then? @Orion1213

OpenStudy (anonymous):

is it \[a_n=a_{n-1}+6~~?\]

OpenStudy (anonymous):

well what does a1 mean? is it the first given term or the first number i have to plug in?

OpenStudy (anonymous):

well if \(a_1=7\) then \(a_0=7-6=1\)... moving forward...\[a_2=7+6=13\]\[a_3=13+6=19\]and so on...

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

i understand how to get the rest of the numbers i was just confused about whether or not 7 is the first term

OpenStudy (anonymous):

but I would choose the last choice.

OpenStudy (anonymous):

7 is really the 1st term noted as \(a_1\)

OpenStudy (anonymous):

yes that is the correct choice... the last choice...

OpenStudy (anonymous):

thank you! do you think you'd be able to help me with another one similar to this? @Orion1213

OpenStudy (anonymous):

sure i still have time...

OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 4, an = 2 • an-1

OpenStudy (anonymous):

so the first term would be 4? and the second one would be found how?

OpenStudy (anonymous):

if \(a_1=4\) then \(a_2=2 \times 4=8\)... moving further...\[a_3=2 \times a_2=2\times8=16\]\[a_4=2 \times a_3=2\times16=32\]and so on...

OpenStudy (anonymous):

got the pattern?

OpenStudy (anonymous):

yes thank you so much!!

OpenStudy (anonymous):

yw... :)

OpenStudy (anonymous):

For this question did I write the equation correctly? A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 200 centimeters tall. an= 200cm+0.5cm(n-1)

OpenStudy (anonymous):

\(a_{n-1}\) refers to the previous term in the sequence...

OpenStudy (anonymous):

yes you write it correct...

OpenStudy (anonymous):

find \[a _{1+1}\] because you a1...similarly repeat \[a _{1+1+1}\] and cycle goes on... so you will get, \[a _{1}=a _{1}+6\] similarly do for rest of them and you will get 7, 13, 19, 25, 31, 37 Hope it's clear @faariat